OpenStudy (esmiiramoslove):

write an algebraic expression that models the word phrase "four more then a number b". (1 Point)

jimthompson5910 (jim_thompson5910):

what are your thoughts on this one? how far did you get?

OpenStudy (esmiiramoslove):

i dont really undersand this

jimthompson5910 (jim_thompson5910):

"a number b" is just a placeholder for a number let's say b = 5

jimthompson5910 (jim_thompson5910):

`"four more then a number b".` would turn into `"four more then a number 5".` what's four more than 5?

OpenStudy (esmiiramoslove):

my choices are A. 4+b B. 4b C. 4-b D. b-4

OpenStudy (esmiiramoslove):

9?

jimthompson5910 (jim_thompson5910):

how did you go from 5 to 9?

jimthompson5910 (jim_thompson5910):

what math operation did you do?

OpenStudy (esmiiramoslove):

add 4

jimthompson5910 (jim_thompson5910):

yes, so whatever the number is, you add 4

jimthompson5910 (jim_thompson5910):

5+4 = 9 b+4 would stay as b+4

jimthompson5910 (jim_thompson5910):

b+4 is the same as 4+b because you can add numbers in any order eg: 2+3 = 3+2 = 5

OpenStudy (esmiiramoslove):

so it would be A?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (esmiiramoslove):

could you help me with more?

jimthompson5910 (jim_thompson5910):

sure, just post a new thread after closing this one